Circular No. 7323

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA AQUILAE 1999 No. 2 Alfredo Pereira, Cabo da Roca, Portugal, reports his visual discovery

of an apparent nova during a search program with 14×100 binoculars on Dec. 1.785 UT at m_{v} ∼ 6.0. D. di Cicco, Sudbury, MA, reports the following precise position from unfiltered CCD frames taken around Dec. 1.95: α = 19^{h}23^{m}05^{s}.38, δ = +4^{o}57′20′′.1 (equinox 2000.0; GSC reference stars), mag 5.2. The Digital Sky Survey reveals a star that has position end figures 05^{s}.32, 19′′.6 (red mag 15.6, blue mag 17.4) in the USNO A2.0 catalogue. Additional visual magnitude estimates: Dec. 1.813, 5.8 (Pereira); 1.819, 5.9 (C. Vitorino, Cabo da Roca, Portugal); 1.842, 5.8 (Vitorino); 1.843, 5.8 (Pereira); 1.939, 5.1 (D. W. E. Green, Cambridge, MA).

HD 209458 E. Poretti, Brera Astronomical Observatory, writes: “Forty-six V pho-

toelectric measurements of HD 209458 obtained with the Merate 0.5-m telescope show that a transit (depth 0.016 ± 0.002 mag) started at Nov. 25.815 UT. Observations began at Nov. 25.74, but stopped at Nov. 25.91 due to large airmass, before egress. Analysis of B photometry is in progress.”

Independent reports have been received from S. Söderhjelm, Lund Observatory; and from N. Robichon and F. Arenou, DASGAL, Observatoire de Paris, Meudon, of the existence of five HD 209458 transits in 89 separate Hipparcos photometric measurements obtained between 1991 Apr. 17 and Nov. 4. Robichon and Arenou report that the mean depth of the transits is 0.025 ± 0.004 magnitude in the Hipparcos H_{p} band, corresponding to a relative depth in the flux of 2.3 ± 0.4 percent (results submitted to A.Ap.). Using the epoch T_{c} of Charbonneau et al. (submitted to Ap.J. Lett.; see also IAUC 7315), periods of 3.524738 ± 0.000015 (Robichon and Arenou) and 3.52473 ± 0.00005 (Söderhjelm) days have been derived. Söderhjelm’s details will appear on IBVS 4816.

COMET P/1999 V1 (CATALINA) Orbital elements (MPC 36654) show this to be a short-period comet:

T = 1999 Oct. 22.1860 TT  ω = 185.^{o}9914
e = 0.588712  Ω = 294.4267 }2000.0
q = 2.961381 AU  i = 15.5816

a = 7.200261 AU n^{o} = 0.0510131 P = 19.32 years

1999 December 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green